# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v1

@@ 4.1.3 Model Robustness Across Different Scales
Table 2: Relative PASS@1 degradation aggregated over CRUX and LCB under direct inference. Weighted Average column uses weights: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC.
@@
We further analyze robustness across different model scales and versions, comparing relative PASS@1 degradation under representative perturbations in Table 2. The weighting details and alternative evaluation schemes are in Appendix C.2.
We observe a clear trend of increasing robustness with model scale in most families. Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B). Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B-Ins) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B-Ins), respectively.
Key Finding: Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
@@ 4.2.1 Effect of CoT Prompting
Table 3: Comparison of relative PASS@1 degradation aggregated over CRUX and LCB for output prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
@@
We first investigate how CoT enhances LLM robustness by code-level step-by-step execution and present the results in Table 3. It effectively enhances model robustness, with average improvements of $10.01$pp for ALL, $9.23$pp for MDC, $4.04$pp for MPS, and a notable $13.57$pp for MHC. It consistently reduces average performance degradation from approximately $24\%$ to $14\%$ on ALL and MDC, indicating that step-by-step reasoning enables LLMs to prioritize executable code semantics over irrelevant segments and cues. The most significant improvement on MHC suggests CoT enhances code reasoning and execution, mitigating over-reliance on misleading hints observed in direct inference in LCB4.1.2). These benefits are amplified in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), which exhibit greater reasoning capacity and noise resistance.
However, CoT does not fully eliminate the model’s reliance on misleading NL information. Performance degradation remains substantial under ALL, MDC, and MPS perturbations, especially for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1). These results highlight a limitation in distinguishing non-functional code segments and misleading information from the program. Interestingly, models tend to incorporate print statements directly into their reasoning, resulting in higher error rates on MPS. We provide detailed case studies for each perturbation in Appendix D (Case 1 to 3) to investigate their inference and illustrate their failure reasons.
Key Finding: CoT prompting significantly enhances LLMs’ robustness by improving reasoning ability and allowing step-by-step execution, especially for the high-capacity models. However, it does not fully mitigate reliance on misleading information and remains constrained by the model’s underlying capabilities.
@@ 4.2.2 Powerful Reasoning and Potential Vulnerability of LRMs
Table 4: Absolute PASS@1 degradation and reasoning token usage (Avg/Max) of LRMs on output prediction across CRUX and LCB (N=1).
@@
We then evaluate the robustness and reasoning behaviors of three LRMs on output prediction tasks, with results summarized in Table 4. All LRMs exhibit remarkable performance and robustness, achieving over 90 PASS@1 accuracy in the VAN setting and having minor performance degradation under the ALL and MPS perturbations. These results suggest that internal reasoning helps models accurately identify and distinguish code segments that are functionally critical for execution, thereby outperforming all generic models with CoT prompting. Moreover, LRMs exhibit consistently stronger robustness against MPS than MDC, indicating that while they can correctly recognize and ignore misleading messages from print statements, they still inherently treat comments as valid execution guidance. This behavior reveals the stable cognitive preference toward specific injection formats, as observed in §3.3.2.
However, this residual cognitive bias leads to a critical vulnerability under the MHC perturbation. Concise and plausible but incorrect output hints can trigger severe reasoning failures in LRMs. We observed that: (1) All LRMs exhibit substantial increases in reasoning token usages under MHC (doubling the token usages in VAN). (2) In extreme cases, QwQ-32B generates over 32,000 tokens in a single task, caused by uncontrolled self-reflection and reasoning collapse (see §5.3 and Appendix D.5). (3) Excessive reasoning leads to more significant degradations in o3-Mini with higher reasoning effort. These observations reveal that logically plausible yet incorrect hints can trigger catastrophic overthinking behavior in LRMs, resulting in significant computational overhead, excessive self-reflection, and severe performance drop.
Key Finding: LRMs exhibit exceptional robustness under ALL and MPS, benefiting from strong internal reasoning capabilities. However, they retain a residual tendency to trust and analyze comments; thereby, plausible but incorrect information can trigger catastrophic overthinking, excessive self-verification, and severe performance degradation, highlighting a critical vulnerability of the internal reasoning mechanism in current LRMs.
@@ 4.3 Impact on Input Prediction
Table 5: Comparison of relative PASS@1 degradation under CRUX for input prediction tasks in direct inference (N=3) and CoT inference (N=1) [TO ABSOLUTE SCORE].
@@
In this section, we analyze the impact of perturbations on input prediction tasks. Consistent with findings from the output prediction in §4.1, the ALL and MHC perturbations cause substantial performance degradation, with relative average drops of $26.63\%$ and $20.04\%$, respectively. These results align with Gu et al. (2024)’s hypothesis that robustness in related coding tasks is correlated. Furthermore, CoT prompting improves model performance by $6.97$pp under ALL and $8.64$pp under MHC, which is less pronounced than output prediction tasks due to the challenges and complexity of the inverse nature of the reasoning process.
Surprisingly, in contrast to output prediction, misleading message perturbations (MDC and MPS) do not degrade model performance in input prediction. Since input prediction allows multiple valid solutions, models can ignore certain branches of code logic, producing reasoning shortcuts. Our misleading messages are intentionally designed to confuse or distract models, but these comments inadvertently reinforce such shortcuts, simplifying models’ reasoning processes. As a result, models tend to produce trivial but valid inputs that bypass critical branches of the program without truly understanding and analyzing the underlying code semantics, which does not reflect genuine robustness in code comprehension. We provide a detailed case study in Appendix D.4 illustrating how models adopt these misleading messages to shortcut their reasoning.
Key Finding: ALL and MHC perturbations significantly degrade model performance, consistent with output prediction results, and reveal robustness challenges in code comprehension. Misleading NL messages accidentally reinforce shortcuts, encouraging models to produce trivial but valid answers without deep understanding.
@@ 5.1 Comment Density and Robustness Degradation
We design a textual density stress test to verify the validity of misleading comment perturbations and evaluate the resilience of different LLMs under varying densities of injected distractions on CRUX dataset. Specifically, we analyze how model performance changes as we gradually increase the comment injection probability $p\in\{0.0,0.1,\ldots,1.0\}$. We present the relative PASS@1 degradation of Qwen models across different perturbation densities in Figure 2. We further investigate the role of reasoning by evaluating CoT inference using Qwen2.5-72B-Ins and the internal reasoning model QwQ-32B, showcasing their robustness trends and reasoning token usages in Table 6.
Despite mild non-linearities caused by the randomness in injection positions and comment content, all models exhibit a clear performance degradation trend as the injection probability $p$ increases. Specifically, smaller models degrade more rapidly, while larger models demonstrate greater resilience under higher densities. This trend extends to CoT prompting and LRMs, with reasoning-enhanced models showing superior robustness against misleading textual distractions. These results validate MDC as a scalable stress-testing technique and an effective method for evaluating the robustness of LLMs under textual perturbations.
While reasoning-enhanced models have similar trends in the stress test, they react differently to dense comment noise. CoT prompting improves robustness by guiding models to selectively focus on the core logic path and ignore distracting signals. However, it can lead to reasoning shortcuts or premature conclusions without sufficiently analyzing relevant information, so token usage is slightly reduced as models do not analyze the comments but reduce the inference effort. In contrast, QwQ-32B processes all available information, including noisy or conflicting cues, and attempts to identify, reconcile, and resolve contradictions before making a final decision. This exhaustive reasoning enables higher vanilla performance and robustness compared to generic models. However, it also risks overthinking and collapse under contradiction-heavy settings, as we discussed in § 4.2.2.
@@ 5.2 Prompting to Ignore Misleading Comments
Table 7: Effect of Comment-Ignoring Prompt (PASS@1).
@@
We conduct a controlled experiment by adding the instruction: “Please ignore the comments and any other non-code elements in the code snippet.” to the prompt on three strong models under direct and CoT inference. From Table 7, even when explicitly instructed to ignore the comments, the PASS@1 scores under direct inference remain similar to those without the instruction, indicating that LLMs still incorporate NL cues into their inference, which is consistent with our finding on their inherent reliance on embedded NL messages in § 4.1.2. Although CoT prompting provides improvements under comment-ignore instructions, the PASS@1 performance of MDC across all models is still far away from the vanilla performance, indicating that even with step-by-step reasoning, LLMs are unable to fully detach from misleading NL distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on surface cues.
5.3 Revealing the Vulnerability of LRMs
Figure 3: Cumulative confusion word trend in QwQ’s output (LCB sample 259 under MHC).
In §4.2.2, we highlighted a critical vulnerability in LRMs. We observed an unhealthy self-reflection loop and a progressive loss of confidence by analyzing a failure case where QwQ-32B generated over 32,000 tokens in response to an MHC-perturbed code.
We manually reviewed and segmented the reasoning process into six stages (see Figure 4 and Appendix D.5). Initially, the model reasons correctly and computes the correct output. However, our plausible yet incorrect hint triggers it to doubt its logic and repeatedly re-execute the same reasoning steps. This recursive self-verification gradually becomes brute-force enumeration, causing the model to hallucinate explanations. Ultimately, the reasoning process collapses, and the model repeatedly outputs over 12,000 “Hmm.”
To further investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words, such as “hmm”, “wait”, “but”, “perhaps” and “mistake”, and visualize the trend in Figure 3. The trend exhibits a striking quadratic growth pattern ($R^{2}=0.9991$), indicating compounding confusion as reasoning progresses. The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, culminating in complete failure and collapse.
@@ 6 Conclusion
Appendix A Prompts
Our prompt instructions for LLMs follow the format designed by CRUX, which is also adopted by LCB, utilizing a two-shot prompt for direct inference and a single-shot prompt for CoT inference to ensure models can follow the instruction and output format effectively. For LRMs, internal reasoning is inherently required before responding. Therefore, applying CoT prompts may lead to redundant reasoning and unfair comparison with standard LLMs under CoT prompting. Considering the fairness, we use direct inference prompts for RM experiments in §4.2.2.
We provide the complete prompts used for output prediction in both direct and CoT inference (Code 1 and 2) and for input prediction in both direct and CoT inference (Code 3 and 4).
A.1 Output Prediction (Direct)
Listing 1: Output Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Ensure the provided expression syntax is correct!
Example 1:
Given the code snippet:
‘‘‘python
def f(n):
return n
and the function call with input arguments:
‘‘‘python
f(17)
Example 2:
Given the code snippet:
‘‘‘python
def f(s):
return s + "a"
and the function call with input arguments:
‘‘‘python
f("x9j")
A.2 Output Prediction (CoT)
Listing 2: Output Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and the function call with input arguments:
‘‘‘python
{input}
Ensure the provided expression syntax is correct!
For example:
Given the code snippet:
‘‘‘python
def f(s):
s = s + s
return "b" + s + "a"
and the input arguments:
‘‘‘python
f("hi")
[THOUGHT]
Lets execute the code step by step:
1. The function f is defined, which takes a single argument s.
5. The return value of the function is therefore "bhihia".
[/THOUGHT]
A.3 Input Prediction (Direct)
Listing 3: Input Prediction Prompt with Direct Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
Example 1:
Given the code snippet:
‘‘‘python
def f(my_list):
count = 0
for i in my_list:
if len(i) % 2 == 0:
count += 1
return count
and output value:
‘‘‘python
3
Example 2:
Given the code snippet:
‘‘‘python
def f(s1, s2):
return s1 + s2
and output value:
‘‘‘python
"banana"
A.4 Input Prediction (CoT)
Listing 4: Input Prediction Prompt with CoT-based Inference.
Given the code snippet:
‘‘‘python
{code}
and output value:
‘‘‘python
{output}
For example:
Given the code snippet:
‘‘‘python
def f(x):
return x + 1
and the output value:
‘‘‘python
17
[THOUGHT]
[/THOUGHT]
Appendix B Illustrative Examples of Perturbation Strategies
In section 3.3, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example. Code 5 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
Listing 5: Vanilla example from LCB (sample 36) for output prediction.
@@ B.1 Renaming Entities
Listing 6: REN-perturbed example from LCB (sample 36) for output prediction.
@@
B.2 Reformatting Conditional Expression
@@
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 7 presents an example of this perturbation.
Listing 7: RTF-perturbed example from LCB (sample 36) for output prediction.
@@
B.3 Inserting Garbage Code

### v2

@@ 4.2.3 Model Robustness Across Different Scales
We further analyze reasoning reliability across different model scales and versions, comparing relative Pass@1 degradation under representative perturbations in Table 2. Due to length constraints, the table, weighting scheme, and additional evaluation setups are provided in Appendix D.2. We observe a clear and consistent trend except for Gemini families. Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models, such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B). Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively.
Table 2: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Absolute Version]
@@ 4.3.1 Effect of CoT Prompting
Table 3: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
@@
We examine how CoT prompting improves LLM robustness through explicit step-by-step execution. From Table 3, we observe that CoT effectively enhances model robustness, with relative improvements of $+10.01$pp for ALL, $+9.23$pp for MDC, $+4.04$pp for MPS, and $+13.57$pp for MHC. The results indicate that step-by-step reasoning guides LLMs to prioritize executable semantics over irrelevant code segments and textual cues. Notably, the substantial improvement in MHC suggests CoT effectively mitigates the over-reliance on reasoning-level hints by re-centering reasoning around executable semantics. Such benefits are particularly evident in high-capacity models (e.g., GPT-4o, Claude-3.5-Sonnet), reflecting enhanced reasoning capabilities and greater noise resilience.
However, CoT prompting cannot entirely eliminate the model’s sensitivity to misleading NL cues and structural distractions. Substantial performance degradation remains under ALL, MDC, and MPS perturbations, particularly for smaller open-sourced models (e.g., Qwen2.5 and LLaMA-3.1). These results underscore an intrinsic limitation of current models in distinguishing non-functional code and misleading information from essential logic. Interestingly, models usually incorporate print statements directly into their reasoning paths, leading to greater degradation under MPS compared to MDC due to self-induced misinterpretation. Since CoT prompting exposes intermediate reasoning processes, we provide detailed case studies for each perturbation in Appendix E (Case 1 to 4) to investigate model inference and illustrate their failure reasons.
Key Finding: CoT prompting significantly improves LLM robustness by enforcing explicit step-by-step reasoning, particularly benefiting larger models. Nevertheless, it cannot fully eliminate models’ reliance on misleading cues, indicating fundamental limitations in reasoning capabilities.
@@ 4.3.2 Powerful Reasoning and Potential Vulnerability of LRMs
We then evaluate the reliability of three LRMs on output prediction tasks, with results summarized in Table 4. Compared to LLMs, LRMs show outstanding performance and robustness, surpassing $90$ Pass@1 in vanilla settings and maintaining minimal performance drop under ALL and MPS perturbations. These results indicate that internal reasoning outperforms CoT prompting in accurately distinguishing essential executable code segments from irrelevant code and misleading distractions.
Notably, LRMs consistently exhibit stronger resistance to MPS compared to MDC, suggesting an inherent cognitive bias toward treating comments as authoritative execution guidance while appropriately disregarding print content. However, this residual cognitive bias exposes LRMs to critical vulnerabilities under reasoning-level perturbation that plausible but incorrect hints can trigger severe reasoning failures. Specifically, we observe the following outcomes under MHC: (i) substantial increases in reasoning token usage across all LRMs, often doubling when compared to VAN; (ii) extreme cases in QwQ-32B—occurring in $4/479$ tasks; see Appendix E.6—where it generates over 32,000 tokens due to uncontrolled recursive self-verification and reasoning collapse; (iii) increased reasoning effort and token usage instead exacerbate performance drops in o3-Mini, suggesting that internal reasoning can amplify confusion when led by contradiction. These findings reveal that reasoning-level contradictories can evoke severe internal reasoning instability, resulting in significant computational overhead, excessive self-reflection, and even severe performance drops.
Table 4: Absolute Pass@1 degradation and reasoning token usage (Avg. and Max.) of LRMs on output prediction across Crux and LCB ($N=1$).
@@
Key Finding: LRMs leverage strong internal reasoning capabilities to achieve remarkable performance and robustness. However, their cognitive tendency toward comments introduces severe vulnerabilities, resulting in doubled token usage and reasoning collapse under MHC perturbation.
@@ 4.4 Impact on Input Prediction
Table 5: Comparison of relative Pass@1 degradation under Crux for input prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
@@
Finally, we analyze how perturbations affect LLMs in input prediction; the results are summarized in Table 5. Consistent with our output prediction findings in §4.2, ALL and MHC perturbations substantially degrade model performance, causing relative average drops of $26.63\%$ and $20.04\%$, respectively. These results support Gu et al. (2024)’s hypothesis that robustness in related coding tasks is interrelated. CoT prompting also improves model performance by $+6.97$pp under ALL and $+8.64$pp under MHC. However, improvements in input prediction are less effective than in output prediction due to the unique challenges of the inverse nature of the reasoning process.
Surprisingly, in contrast to output prediction, contextual-level textual perturbations (MDC and MPS) do not negatively impact LLMs in input prediction tasks. Since input prediction tasks accept multiple valid solutions, models can bypass reasoning about certain code branches for a correct answer. Our misleading NL messages, initially designed to confuse and distract models, inadvertently reinforce such shortcuts, simplifying models’ reasoning processes. Consequently, these perturbations unintentionally provide shallow reasoning strategies, so those improvements do not reflect the genuine robustness of LLMs in code comprehension. We provide a detailed case study in Appendix E.5, illustrating how models adopt misleading messages to shortcut their reasoning.
Key Finding: ALL and MHC significantly degrade performance on input prediction, highlighting fundamental reasoning limitations. Conversely, misleading messages accidentally simplify reasoning, prompting models to produce superficial but valid answers without true understanding.
@@ 5.1 Comment Density and Robustness Degradation
We design a textual density stress test to confirm the validity of MDC perturbations and evaluate model robustness under varying distraction densities on Crux dataset. Specifically, we examine how relative Pass@1 drop across different Qwen models as we gradually increase the comment injection probability $p\in\{0.0,0.1,\dots,1.0\}$, illustrated in Figure 1. We further analyze the role of reasoning by evaluating CoT inference using Qwen2.5-72B and internal reasoning using QwQ-32B, with detailed results on token usage summarized in Table 6.
All models exhibit a clear performance degradation trend as injection probability $p$ increases, despite minor non-linearities due to randomness in insertion positions and comment content. Notably, smaller models degrade more rapidly, while larger models demonstrate greater resilience at high densities. This trend persists in reasoning-enhanced models (CoT and LRMs), showing superior robustness against dense textual distractions. These results confirm the validity of injecting misleading messages (MDC and MPS) as a scalable stress-testing technique and an effective strategy to evaluate LLM robustness under contextual-level perturbations.
While Qwen2.5-72B with CoT prompting and QwQ-32B have similar trends in the stress test, they respond and reason differently under dense comment noise. CoT prompting improves model robustness by guiding models to selectively focus on the essential code logic and ignore distracting messages. However, it can lead to premature conclusions without sufficiently analyzing the code, so token usage is slightly reduced as models do not analyze the comments and reduce the inference effort. In contrast, QwQ-32B processes all available information and attempts an exhaustive analysis, actively resolving contradictions before making a final decision. While this comprehensive analysis achieves higher vanilla performance and overall robustness, it risks severe reasoning collapse under heavily contradictory conditions, as discussed in §4.3.2.
@@ 5.2 Prompting to Ignore Misleading Comments
Table 7: Impact of explicit comment-ignoring prompts under MDC and MHC perturbations.
@@
Next, we conduct a controlled experiment by instructing models: “Please ignore the comments and any other non-code elements in the code snippet.” Table 7 summarizes Pass@1 results on LCB for three high-performing models under both direct and CoT inference modes. Despite explicit instructions to ignore comments, the Pass@1 accuracy remains largely unaffected under direct inference, supporting our findings from §4.2.1 about LLMs’ inherent reliance on embedded NL cues. Although we observe improvements after employing CoT prompting, the Pass@1 performance under MDC is still substantially below vanilla settings. These results indicate that even with step-by-step execution, LLMs cannot fully detach from misleading textual distractions, highlighting a fundamental limitation in their reasoning controllability and dependence on NL cues.
5.3 Revealing the Vulnerability of Self-reflection Mechanisms
Figure 2: Quadratic increase ($R^{2}=0.9991$) in confusion tokens (e.g., “hmm”, “wait”, “perhaps”) during QwQ-32B’s reasoning process on LCB sample 259 under the MHC perturbation.
In §4.3.2, we identified a critical vulnerability in QwQ-32B’s internal reasoning. By analyzing failure cases that generated over 32,000 tokens under the MHC perturbation (occurring in 4 out of 479 tasks), we observed an unhealthy self-reflection loop and a progressive loss of confidence. We manually segmented the reasoning process into six stages (see Figure 4 and Appendix E.6). Initially, the model accurately computes the correct output, but the injected incorrect hint triggers it to repeatedly doubt its logic and re-execute the reasoning steps. This recursive verification gradually evolves into a brute-force enumeration, causing the model to generate hallucinated explanations. Ultimately, the reasoning process entirely collapses, with the model repeatedly generating over 12,000 “Hmm.”
To investigate the model’s confidence before collapsing, we track the cumulative frequency of confusing words and visualize the trend in Figure 2. The trend exhibits a striking quadratic growth pattern, indicating compounding confusion as reasoning progresses. The quadratic trajectory supports our stage-wise analysis: the model initially shows stable reasoning, but once the misleading hint is triggered, confusion accumulates, confidence undermines, culminating in complete failure.
@@ 6 Conclusion
Appendix A Illustrative Examples of Perturbation Strategies
Figure 3: Overview of CodeCrash pipeline. Code samples are first standardized into a clean vanilla version. Perturbations are visualized with yellow highlights. Each perturbed code is validated and executed before passing to LLMs. [Back To Execution Pipeline]
In section 3.3, we briefly describe all perturbation strategies employed in this paper. In this appendix section, we illustrate the detailed implementation for all perturbations with a LCB sample 36 as the example. Code 1 presents the base vanilla example where the formal answer is minimumCost(s = ’0011’) == 2.
Code 1: Vanilla example from LCB (sample 36) for output prediction.
@@ A.1 Renaming Entities
Code 2: REN-perturbed example from LCB (sample 36) for output prediction.
@@
A.2 Reformatting Conditional Expression
@@
Here, {cond} denotes the original condition, and {para} is a randomly selected valid parameter or variable from the current function scope. Code 3 presents an example of this perturbation.
Code 3: RTF-perturbed example from LCB (sample 36) for output prediction.
@@
A.3 Inserting Garbage Code